# RefAM: Attention Magnets for Zero-Shot Referral Segmentation

### v1

@@ Abstract
Most existing approaches to referring segmentation achieve strong performance only through fine-tuning or by composing multiple pre-trained models, often at the cost of additional training and architectural modifications. Meanwhile, large-scale generative diffusion models encode rich semantic information, making them attractive as general-purpose feature extractors. In this work, we introduce a new method that directly exploits features, attention scores, from diffusion transformers for downstream tasks, requiring neither architectural modifications nor additional training. To systematically evaluate these features, we extend benchmarks with vision-language grounding tasks spanning both images and videos. Our key insight is that stop words act as attention magnets: they accumulate surplus attention and can be filtered to reduce noise. Moreover, we identify global attention sinks (GAS) emerging in deeper layers and show that they can be safely suppressed or redirected onto auxiliary tokens, leading to sharper and more accurate grounding maps. We further propose an attention redistribution strategy, where appended stop words partition background activations into smaller clusters, yielding sharper and more localized heatmaps. Building on these findings, we develop RefAM, a simple training-free grounding framework that combines cross-attention maps, GAS handling, and redistribution. Across zero-shot referring image and video segmentation benchmarks, our approach consistently outperforms prior methods, establishing a new state of the art without fine-tuning or additional components.
@@ 1 Introduction
Diffusion transformers (DiTs) have rapidly advanced generative modeling and, more recently, been adopted as powerful feature extractors for downstream vision–language tasks such referring object segmentation (Ni et al., 2023). Their cross-attention maps encode rich spatial and semantic information without task-specific training, making them attractive for training-free and zero-shot applications. However, attention in transformers is also known to exhibit emergent behaviors that are not always semantically meaningful. In large language models, for instance, certain tokens, often first tokens, attract disproportionately high attention while carrying little to no semantic content, a phenomenon referred to as attention sinks or massive activations (Xiao et al., 2024; Yona et al., 2025; Jin et al., 2025; Sun et al., 2024a; Barbero et al., 2025).
In this paper, we extend this observation to generative diffusion transformers and show that they exhibit similar attention sink behaviors when applied to vision–language grounding tasks. Specifically, we uncover language–vision attention sinks, where stop words emerge as high-attention tokens despite lacking semantic value. We find two distinct patterns. First, a small set of stop words consistently act as global attention sinks (GASs) in the later layers of DiTs: they attend almost uniformly across text and image tokens, and filtering their channels does not harm downstream performance. Second, other stop words behave as local background attractors, drawing attention toward irrelevant regions. Surprisingly, appending additional stop words introduces more such attractors, which redistributes background attention and yields cleaner heatmaps. We further show that replacing stop words with random vectors also improves results, but real stop words are consistently more effective, likely due to their repeated presence during pretraining.
These findings suggest that stop words can serve as a simple yet effective tool for attention redistribution. Building on this, we propose RefAM, a training-free grounding method, that augments referring expressions with stop words, filters their attention maps, and aggregates the remaining cross-attention for grounding. This approach requires no modifications to the diffusion model, no additional supervision, and generalizes to both image and video tasks.
@@
  • We introduce a stop-word based attention redistribution strategy, where auxiliary stop words act as magnets that absorb surplus background attention, yielding cleaner and more localized cross-attention maps.
  • We achieve state-of-the-art zero-shot referring segmentation on both image and video benchmarks using RefAM, features from diffusion transformers, outperforming prior training-free methods without fine-tuning or auxiliary components.
@@ 2 Related Work
High-Norm Tokens Across Transformer Architectures. Recent research has identified tokens exhibiting high-norm activations across various domains, including language models (Xiao et al., 2024; Yona et al., 2025; Jin et al., 2025; Sun et al., 2024a; Barbero et al., 2025), vision models (Kang et al., 2025; Darcet et al., 2024; Jiang et al., 2025; Wang et al., 2024), and vision-language models (An et al., 2025; Woo et al., 2025). In language models, these tokens are referred to as attention sinks (Xiao et al., 2024; Yona et al., 2025; Barbero et al., 2025) or massive activations (Jin et al., 2025; Sun et al., 2024a). In vision models, similar phenomena are termed registers (Darcet et al., 2024; Jiang et al., 2025), visual attention sinks (Kang et al., 2025), or defective path tokens (Wang et al., 2024). In vision-language models, this phenomenon has been described as attention deficiency (An et al., 2025) or blind tokens (Woo et al., 2025), reporting individual visual tokens that consistently receive disproportionately high attention. These studies consistently show that a small fraction of tokens absorb disproportionately high attention, often without semantic relevance. In our work, we identify global attention sinks (GAS), tokens that span both language and visual streams under the same query and systematically suppress useful signals in both modalities, simultaneously.
@@
A common explanation attributes this behavior to the softmax normalization constraint: attention weights must sum to one, even when the query lacks a strong contextual match (Xiao et al., 2024). In such cases, attention is distributed toward tokens that act as “sinks”. Other studies offer complementary views: Jin et al.(Jin et al., 2025) point to positional encodings; Sun et al.(Sun et al., 2024a) implicate learned bias terms; and Wang et al. (Wang et al., 2024) connect the effect to the power method, where repeated matrix multiplications amplify dominant directions in feature space. While many approaches aim to mitigate this behavior, we instead draw on it: our method introduces attention magnets, tokens, e.g., stop words, that deliberately absorb surplus attention to help redistribute focus more effectively in generative models.
@@
Referring object segmentation. Referring object segmentation aims to localize a region in an image or video based on a natural language expression. We address both static and temporal settings. (1) Referring Image Object Segmentation (RIOS). Traditional methods are supervised (Kazemzadeh et al., 2014; Ding et al., 2020; Feng et al., 2021; Li et al., 2018), but recent zero-shot approaches leverage pre-trained models. Global-Local (Yu et al., 2023) extracts CLIP-based features from mask proposals, and Ref-Diff (Ni et al., 2023) uses diffusion priors. HybridGL (Liu & Li, 2025) fuses global-local context with spatial cues. These methods first generate multiple object proposals and then score masks by their similarity to text embeddings. (2) Referring Video Object Segmentation (RVOS). Temporal methods like LoSh (Yuan et al., 2024) and WRVOS (Zhao et al., 2023) require supervision. In contrast, recent zero-shot methods such AL-Ref-SAM (Ren et al., 2024) use pre-trained image grounding GroundedSAM (Ren et al., 2024) and SAM2 (Ravi et al., 2025) models for frame-wise grounding with minimal adaptation. Unlike prior approaches, our method unifies both tasks under a single framework using diffusion-derived features, operating in a training-free, zero-shot setting.
@@ 3 Semantic Features from Diffusion Transformers
We describe how semantic features are extracted from rectified-flow diffusion transformers (DiTs) and how stop words, acting as attention magnets, enable robust referral segmentation. Our method consists of three components: (i) extracting cross-attention maps from DiTs, (ii) identifying and filtering attention sinks, and (iii) redistributing surplus attention through additional tokens (Figure 2).
@@ 3.1 Feature Extraction from DiTs
Rectified flow models (Labs, 2024; Team, 2024) combine an image-to-latent encoder–decoder with a DiT (Peebles & Xie, 2023) backbone. The encoder compresses inputs into a latent space, while the DiT performs denoising through a sequence of transformer blocks. Architectures may interleave double-stream blocks, which process text and visual tokens separately before merging in attention, and single-stream blocks, which operate on concatenated tokens with shared weights.
Given a clean latent $X_{0}$, the rectified flow forward process perturbs it as
$X_{t}=(1-\sigma)X_{0}+\sigma\epsilon,\quad\epsilon\sim\mathcal{N}(0,I).$
While the DiT is trained to predict the noise $\epsilon$, its intermediate activations capture rich semantic information. In particular, cross-attention maps between text and image tokens provide spatial grounding signals. For the denoising process itself, the model uses either the source prompt (if available) or an empty prompt. In parallel, we collect features from a separate text branch that encodes the referring expression, similarly as in (Helbling et al., 2025). This branch is used exclusively for feature extraction and has no effect on the denoising trajectory. Unlike prior work that primarily relies on U-Net features (Tang et al., 2023a; Zhang et al., 2023), we exploit these transformer attention maps directly, which we find more effective for referring segmentation.
@@ 3.3 Emergence of Semantic Information in DiT
Late layers (emergence of GAS). In later layers, semantic alignment sharpens, but we also consistently observe Global Attention Sinks (GAS). These are tokens, most often stop words, that allocate unusually high and nearly uniform attention across both text and image tokens (Fig. 1). We identify GAS by computing per-token text-to-text activations and marking tokens whose average mass is $10\times$ higher than the mean across all layers and all tokens. In Fig. 1, we visualize layer-averaged text-to-text and text-to-visual attention for each token. On the left, tokens #1 (‘_a’) and #16 (‘<</s>>’) exhibit uniformly high attention across both textual and visual tokens, characteristic of global attention sinks. Typically, 1–3 GAS tokens appear per sequence.
@@ 3.4 Interpretation of Global Attention Sinks
Next, we analyze the role of Global Attention Sinks (GAS) and their impact on referral segmentation.
@@ 3.5 Redistribution strategy with attention magnets
Practical effect. The combined mechanism, (i) redirecting global sinks into magnets and (ii) partitioning background noise across multiple attractors, consistently improves grounding. Foreground maps become sharper and more concentrated, while meaningful tokens preserve their semantic roles. Crucially, the approach is entirely training-free: it leverages inductive behavior already learned during pretraining (e.g., frequent exposure to stop words) rather than introducing new parameters.
@@ 4 Results
Datasets. We evaluate our method on the standard benchmarks for referring image and video segmentation tasks. For referring image segmentation (RIOS), we use RefCOCO/+/g (Kazemzadeh et al., 2014), containing referring expressions for objects in COCO images (Lin et al., 2014). For referring video segmentation (RVOS), we use Ref-DAVIS17 (Khoreva et al., 2019), Ref-YouTube-VOS (Seo et al., 2020) and MeViS (Ding et al., 2023), which provides video object masks and expressions for sequences. MeViS is a newly established dataset that is targeted at motion information analysis and its test set consists of 50 videos and 793 annotations. The Ref-YouTube-VOS stands out as the most extensive RVOS dataset, comprising 202 videos and 834 annotations. Ref-DAVIS17 builds upon DAVIS17 (Khoreva et al., 2019) and contains 30 videos with 244 annotations.
Implementation Details. As attention magnets, we append “_”, “with”, “to”, “and” stop words and “pink” as and auxiliary color that redistributes some of the meaningful GAS tokens from the referral expression to ours attention magnets. We filter out stop words used not only as attention magnets, but also stop words within the referring expressions, and the end-of-sequence (¡/s¿) token. We use the same preprocessing strategy as HybridGL (Liu & Li, 2025) to extract noun phrases (NP) and spatial bias (SB) from the referring expression using spacy package, ensuring a fair comparison. For referring image object segmentation (RIOS), we use the FLUX model (Labs, 2024) and collect features from timestep 750. For referring video object segmentation (RVOS), we use the Mochi model (Team, 2024) and collect features from timestep 990. To produce final attention map, we aggregate attention maps across all transformer blocks if not stated differently. Since the COCO dataset already provides captions with its annotations, we use these directly to guide feature extraction. We use chatGPT4o to generate captions for DAVIS, Ref-YouTube-VOS and MeViS test videos.
Metric Method Vision Backbone Pre-trained Model RefCOCO RefCOCO+ RefCOCOg
@@
VLM-VG (Wang et al., 2025) R101 $\text{COCO}$$^{*}$, $\text{VLM-VG}$$^{*}$ 45.40 48.00 41.40 37.00 40.70 30.50 42.80 44.10
@@
Grad-CAM (Selvaraju et al., 2017a) R50 SAM, CLIP 23.44 23.91 21.60 26.67 27.20 24.84 23.00 23.91
@@
mIoU weakly-supervised methods
CLRL (Lee et al., 2023) ViT-B - 31.06 32.30 30.11 31.28 32.11 30.13 32.88 -
PPT (Dai & Yang, 2024) ViT-B SAM 46.76 45.33 46.28 45.34 45.84 44.77 42.97 -
zero-shot methods w/ additional training
@@
VLM-VG (Wang et al., 2025) R101 $\text{COCO}$$^{*}$, $\text{VLM-VG}$$^{*}$ 49.90 53.10 46.70 42.70 47.30 36.20 48.00 48.50
@@
Grad-CAM (Selvaraju et al., 2017a) R50 SAM, CLIP 30.22 31.90 27.17 33.96 25.66 32.29 33.05 32.50
@@
Table 2: Comparison with state-of-the-art zero-shot methods on RefCOCO, RefCOCO+, and RefCOCOg. The top two results in each setting (without additional training) are marked in bold and underlined, respectively. * denotes use of extra training data beyond the task-specific set.
4.1 SOTA Comparison
We evaluate RefAM on referral image object segmentation (RIOS) on RefCOCO, RefCOCO+, and RefCOCOg datasets using oIoU and mIoU metrics in Tab. 2 and on referral video object segmentation (RVOS) Ref-DAVIS17, Ref-YouTube-VOS and MeViS datasets using standard $\mathcal{J}\&\mathcal{F}$ metrics in  Tab. 3. Our method significantly outperforms prior training-free approaches on both RIOS and RVOS, establishing new state-of-the-art results across all datasets. Notable RIOS baselines include Ref-Diff (Ni et al., 2023), MaskCLIP (Zhou et al., 2022), Global-Local (Yu et al., 2023), and the recent HybridGL (Liu & Li, 2025), which rely on complex modeling of spatial or relational cues. In contrast, our approach is simple and leverages the semantic structure learned by pretrained generative models. In particular, compared to HybridGL—the strongest prior zero-shot method—RefAM achieves an absolute gain of +2.5 mIoU on RefCOCOg test and +1.8 mIoU on RefCOCO+ testB. On RefCOCO+ testA, RefAM improves mIoU by more than 9 points over Ref-Diff and by over 12 points over Global-Local. Despite relying only on frozen FLUX features and SAM segmentation method, our approach achieves performance competitive with, and in some cases exceeding, methods that incorporate additional task-specific training or fine-tuning. In Tab. 3, RefAM outperforms all prior training-free baselines and narrowing the gap to recent methods such as Grounded-SAM (Kirillov et al., 2023), Grounded-SAM2 (Ren et al., 2024), and AL-Ref-SAM (Huang et al., 2025), which are pretrained with image grounding datasets. These results demonstrate that carefully leveraging diffusion features, without retraining, is sufficient to close the gap with supervised and weakly supervised methods, while maintaining the simplicity and generality of a fully training-free pipeline.

### v2

@@ Abstract
Most existing approaches to referring segmentation achieve strong performance only through fine-tuning or by composing multiple pre-trained models, often at the cost of additional training and architectural modifications. Meanwhile, large-scale generative diffusion models encode rich semantic information, making them attractive as general-purpose feature extractors. In this work, we introduce a new method that directly exploits features, attention scores, from diffusion transformers for downstream tasks, requiring neither architectural modifications nor additional training. To systematically evaluate these features, we extend benchmarks with vision–language grounding tasks spanning both images and videos. Our key insight is that stop words act as attention magnets: they accumulate surplus attention and can be filtered to reduce noise. Moreover, we identify global attention sinks (GAS) emerging in deeper layers and show that they can be safely suppressed or redirected onto auxiliary tokens, leading to sharper and more accurate grounding maps. We further propose an attention redistribution strategy, where appended stop words partition background activations into smaller clusters, yielding sharper and more localized heatmaps. Building on these findings, we develop RefAM, a simple training-free grounding framework that combines cross-attention maps, GAS handling, and redistribution. Across zero-shot referring image and video segmentation benchmarks, our approach achieves strong performance and surpasses prior methods on most datasets, establishing a new state of the art without fine-tuning, additional components and complex reasoning.
@@ 1 Introduction
Diffusion transformers (DiTs) have rapidly advanced generative modeling and, more recently, been adopted as powerful feature extractors for downstream vision–language tasks such as referring object segmentation (Ni et al., 2023). Their cross-attention maps encode rich spatial and semantic information without task-specific training, making them attractive for training-free and zero-shot applications. However, attention in transformers is also known to exhibit emergent behaviors that are not always semantically meaningful. In large language models, for instance, certain tokens, often first tokens, attract disproportionately high attention while carrying little to no semantic content, a phenomenon referred to as attention sinks or massive activations (Xiao et al., 2024; Yona et al., 2025; Jin et al., 2025; Sun et al., 2024a; Barbero et al., 2025).
We extend this observation to generative diffusion transformers and show that they exhibit similar attention sink behaviors when applied to vision–language grounding tasks. Specifically, we uncover language–vision attention sinks, where stop words emerge as high-attention tokens despite lacking semantic value. We find two distinct patterns. First, a small set of stop words consistently act as global attention sinks (GASs) in the later layers of DiTs: they attend almost uniformly across text and image tokens, and filtering their channels does not harm downstream performance. Second, other stop words behave as local background attractors, drawing attention toward irrelevant regions. Surprisingly, appending additional stop words introduces more such attractors, which redistributes background attention and yields cleaner heatmaps. We further show that replacing stop words with random vectors also improves results, but real stop words are consistently more effective, likely due to their repeated presence during pretraining.
These findings suggest that stop words can serve as a simple yet effective tool for attention redistribution. Building on this, we propose RefAM, a training-free grounding method, that augments referring expressions with stop words, filters their attention maps, and aggregates the remaining cross-attention for grounding. This approach requires neither modifications to the diffusion model, nor additional supervision, and generalizes to both image and video tasks.
@@
  • We introduce RefAM, a stop-word based attention redistribution strategy, where added stop words act as magnets that absorb surplus attention and enable cleaner cross-attention maps.
  • We achieve state-of-the-art or competitive results for zero-shot referring segmentation on image and video benchmarks using RefAM, features from diffusion transformers, outperforming prior training-free methods without fine-tuning, auxiliary components or complex reasoning.
We will make all our code public to make our results reproducible for the broader community.
@@ 2 Related Work
High-Norm Tokens Across Transformer Architectures. Recent research has identified tokens exhibiting high-norm activations across various domains, including language models (Xiao et al., 2024; Yona et al., 2025; Jin et al., 2025; Sun et al., 2024a; Barbero et al., 2025), vision models (Kang et al., 2025; Darcet et al., 2024; Jiang et al., 2025; Wang et al., 2024a), and vision-language models (An et al., 2025; Woo et al., 2024). In language models, these tokens are referred to as attention sinks (Xiao et al., 2024; Yona et al., 2025; Barbero et al., 2025) or massive activations (Jin et al., 2025; Sun et al., 2024a). In vision models, similar phenomena are termed registers (Darcet et al., 2024; Jiang et al., 2025), visual attention sinks (Kang et al., 2025), or defective path tokens (Wang et al., 2024a). In vision-language models, this phenomenon has been described as attention deficiency (An et al., 2025) or blind tokens (Woo et al., 2024), reporting individual visual tokens that consistently receive disproportionately high attention. These studies consistently show that a small fraction of tokens absorb disproportionately high attention, often without semantic relevance. In our work, we identify global attention sinks (GAS), tokens that span both language and visual streams under the same query and systematically suppress useful signals in both modalities, simultaneously.
@@
A common explanation attributes this behavior to the softmax normalization constraint: attention weights must sum to one, even when the query lacks a strong contextual match (Xiao et al., 2024). In such cases, attention is distributed toward tokens that act as “sinks”. Other studies offer complementary views: Jin et al.(Jin et al., 2025) point to positional encodings; Sun et al.(Sun et al., 2024a) implicate learned bias terms; and Wang et al. (Wang et al., 2024a) connect the effect to the power method, where repeated matrix multiplications amplify dominant directions in feature space. While many approaches aim to mitigate this behavior, we instead draw on it: our method introduces attention magnets, tokens, e.g., stop words, that deliberately absorb surplus attention to help redistribute focus more effectively in generative models. Unlike Darcet et al. (2024), who require re-training the model with learnable “register” tokens to accumulate global information, our approach is training-free. We identify that multimodal attention sinks emerge naturally in pre-trained DiTs and that existing stop words effectively fulfill this role. Furthermore, while registers are designed to store information, our attention magnets function as garbage collectors, filtering out noise to sharpen the grounding signal.
@@
Referring object segmentation. Referring object segmentation aims to localize a region in an image or video based on a natural language expression. We address both static and temporal settings. (1) Referring Image Object Segmentation (RIOS). Traditional methods are supervised (Kazemzadeh et al., 2014; Ding et al., 2020; Feng et al., 2021; Li et al., 2018), but recent zero-shot approaches leverage pre-trained models. Global-Local (Yu et al., 2023) extracts CLIP-based features from mask proposals, and Ref-Diff (Ni et al., 2023) uses diffusion priors. HybridGL (Liu & Li, 2025) fuses global-local context with spatial cues. These methods first generate multiple object proposals and then score masks by their similarity to text embeddings. (2) Referring Video Object Segmentation (RVOS). Temporal methods like LoSh (Yuan et al., 2024) and WRVOS (Zhao et al., 2023) require supervision. In contrast, recent zero-shot methods such as AL-Ref-SAM (Ren et al., 2024) use pre-trained image grounding GroundedSAM (Ren et al., 2024) and SAM2 (Ravi et al., 2025) models for frame-wise grounding with minimal adaptation. Unlike prior approaches, our method unifies both tasks under a single framework using diffusion-derived features, operating in a training-free, zero-shot setting.
@@ 3 Semantic Features from Diffusion Transformers
We describe how semantic features are extracted from rectified-flow diffusion transformers (DiTs) and how stop words, acting as attention magnets (AM), enable robust referral segmentation. Our method consists of three components: (i) extracting cross-attention maps from DiTs, (ii) identifying and filtering attention sinks, and (iii) redistributing surplus attention through additional tokens (Figure 2).
@@ 3.1 Feature Extraction from DiTs
Rectified flow models (Labs, 2024; Team, 2024) combine an image-to-latent encoder–decoder with a DiT (Peebles & Xie, 2023) backbone. The encoder compresses inputs into a latent space, while the DiT performs denoising through a sequence of transformer blocks. Architectures may interleave double-stream blocks, which process text and visual tokens separately before merging in attention, and single-stream blocks, which operate on concatenated tokens with shared weights. Given a clean latent $X_{0}$, the rectified flow forward process perturbs it as
$X_{t}=(1-\sigma_{t})X_{0}+\sigma_{t}\epsilon,\quad\epsilon\sim\mathcal{N}(0,I).$
While the DiT is trained to predict the noise $\epsilon$, its intermediate activations capture rich semantic information. In particular, cross-attention maps between text and image tokens provide spatial grounding signals. For the denoising process itself, the model uses either the source prompt (if available) or an empty prompt. In parallel, we collect features from a separate text branch that encodes the referring expression, similarly as in (Helbling et al., 2025). This branch is used exclusively for feature extraction and has no effect on the denoising trajectory. Unlike prior work that primarily relies on U-Net features (Tang et al., 2023; Zhang et al., 2023), we exploit these transformer attention maps directly, which we find more effective for referring segmentation.
@@ 3.3 Emergence of Semantic Information in DiT
Late layers (emergence of GAS). In later layers, semantic alignment sharpens, but we also consistently observe Global Attention Sinks (GAS). These are tokens, most often stop words, that allocate unusually high and nearly uniform attention across both text and image tokens (Fig. 1). We identify GAS by computing per-token text-to-text activations and marking tokens whose average mass is $10\times$ higher than the mean across all layers and all tokens. In Fig. 1, we visualize layer-averaged text-to-text and text-to-visual attention for each token. On the left, tokens #1 (_a) and #16 (</s>) exhibit uniformly high attention across both textual and visual tokens, characteristic of global attention sinks. Typically, 1–3 GAS tokens appear per sequence.
@@ 3.4 Interpretation of Global Attention Sinks
We analyze the role of Global Attention Sinks (GAS) and their impact on referral segmentation.
@@ 3.5 Redistribution Strategy with Attention Magnets
Practical effect. The combined mechanism, (i) redirecting global sinks into magnets and (ii) partitioning background noise across multiple attractors, consistently improves grounding. Foreground maps become sharper and more concentrated, while meaningful tokens preserve their semantic roles. Crucially, this is entirely training-free: it leverages inductive behavior already learned during pretraining (e.g., frequent exposure to stop words) rather than introducing new parameters. This strategy is grounded in recent NLP findings where specific tokens (e.g., punctuation or start-of-sentence tokens) act as attention sinks to stabilize inference (Xiao et al., 2024). We observe a similar phenomenon in multimodal DiTs: stop words naturally attract surplus attention mass. By explicitly appending these “magnets”, we provide a designated destination for background noise, preventing it from contaminating semantic tokens.
@@ 4 Results
Datasets. We evaluate our method on the standard benchmarks for referring image and video segmentation tasks. For referring image segmentation (RIOS), we use RefCOCO/+/g (Kazemzadeh et al., 2014), containing referring expressions for objects in COCO images (Lin et al., 2014). For referring video segmentation (RVOS), we use Ref-DAVIS17 (Khoreva et al., 2019), Ref-YouTube-VOS (Seo et al., 2020) and MeViS (Ding et al., 2023), which provides video object masks and expressions for sequences. MeViS is a newly established dataset that is targeted at motion information analysis and its test set consists of 50 videos and 793 annotations. The Ref-YouTube-VOS stands out as the most extensive R-VOS dataset, comprising 202 videos and 834 annotations. Ref-DAVIS17 builds upon DAVIS17 (Khoreva et al., 2019) and contains 30 videos with 244 annotations.
Implementation Details. As attention magnets, we append stop words like “_”, “with”, “to”, “and” and some auxiliary colors like “pink”, “black”, that redistributes some of the meaningful GAS tokens from the referral expression to ours attention magnets. We filter out stop words used not only as attention magnets, but also stop words within the referring expressions, and the end-of-sequence (</s>) token. We utilize spaCy library to extract noun phrases (NP) and spatial bias (SB) from referring expressions. For referring image object segmentation (RIOS), we use the FLUX (Labs, 2024) and SD3.5 (Esser et al., 2024) models and collect features from timestep 520 and 640. For referring video object segmentation (RVOS), we use the Mochi model (Team, 2024) and collect features from timestep 990. To produce final attention map, we aggregate attention maps across all transformer blocks if not stated differently. Since the COCO dataset already provides captions with its annotations, we use these directly to guide feature extraction. We use chatGPT4o to generate captions for DAVIS, Ref-YouTube-VOS and MeViS test videos.
Metric Method Vision Backbone Pre-trained Model CLIP RefCOCO RefCOCO+ RefCOCOg
@@
VLM-VG (Wang et al., 2024b) R101 $\text{COCO}$$^{*}$, $\text{VLM-VG}$$^{*}$ 45.40 48.00 41.40 37.00 40.70 30.50 42.80 44.10
@@
Grad-CAM (Selvaraju et al., 2017) R50 SAM, CLIP 23.44 23.91 21.60 26.67 27.20 24.84 23.00 23.91
@@
[…]
mIoU zero-shot methods w/ additional training
@@
VLM-VG (Wang et al., 2024b) R101 $\text{COCO}$$^{*}$, $\text{VLM-VG}$$^{*}$ 49.90 53.10 46.70 42.70 47.30 36.20 48.00 48.50
@@
Grad-CAM (Selvaraju et al., 2017) R50 SAM, CLIP 30.22 31.90 27.17 33.96 25.66 32.29 33.05 32.50
@@
[…]
Table 2: Comparison with state-of-the-art zero-shot methods on RefCOCO, RefCOCO+, and RefCOCOg. The top three results in each setting (without additional training) are marked in bold, underlined, italicized, respectively. * denotes use of extra training data beyond the task-specific set.
4.1 Quantitative Analysis
We evaluate RefAM on referral image object segmentation (RIOS) on RefCOCO, RefCOCO+, and RefCOCOg datasets using oIoU and mIoU metrics in Sec. 4 and on referral video object segmentation (RVOS) Ref-DAVIS17, Ref-YouTube-VOS and MeViS datasets using standard $\mathcal{J}\&\mathcal{F}$ metrics in  Tab. 3. Our method significantly outperforms prior training-free approaches on both RIOS and RVOS benchmarks, achieving new state-of-the-art results on 4 out of 6 datasets, while remaining competitive on RefCOCO+ and RefCOCOg. Notable RIOS baselines include Ref-Diff (Ni et al., 2023), MaskCLIP (Zhou et al., 2022), Global-Local (Yu et al., 2023), and the recent HybridGL (Liu & Li, 2025), which rely on complex modeling of spatial or relational cues. In contrast, our approach is simple and leverages the semantic structure learned by pretrained generative models. In particular, compared to HybridGL—the strongest prior zero-shot method—RefAM achieves an absolute gain of +4.45 mIoU on RefCOCO testB. On RefCOCOg test, RefAM improves mIoU by more than 4 points over Ref-Diff and by over 8 points over Global-Local. Despite relying only on frozen FLUX features and SAM segmentation method, our approach achieves performance competitive with, and in some cases exceeding, methods that incorporate additional task-specific training or fine-tuning. In Tab. 3, RefAM outperforms all prior training-free baselines and narrowing the gap to recent methods such as Grounded-SAM (Kirillov et al., 2023), Grounded-SAM2, and AL-Ref-SAM (Ren et al., 2024), which are pretrained with image grounding datasets. These results demonstrate that carefully leveraging diffusion features, without retraining, is sufficient to close the gap with supervised and weakly supervised methods, while maintaining the simplicity and generality of a fully training-free pipeline.
Inference Efficiency. While our method utilizes DiT backbones, it avoids the complex auxiliary modules found in prior works. For instance, HybridGL (Liu & Li, 2025) relies on multiple inference passes and proposal networks, resulting in a reported total inference time of $\sim$1.1 seconds per image. In contrast, RefAM requires approximately 460ms per image (using FLUX-dev on an A100 GPU), making it significantly faster than the strongest training-free baselines while achieving higher or competitive accuracy. Memory usage ($\sim$22GB) remains within standard research hardware limits for large-scale foundation models.